# ChunkKV: Semantic-Preserving KV Cache Compression for Efficient Long-Context LLM Inference

### v1

@@ B.3 Needle-In-A-Haystack
Figure 17 and 18 visualizes the performance of ChunkKV on the NIAH benchmark for LLaMA-3-8B-Instruct and Mistral-7B-Instruct with a KV cache size of 128 under 8k and 32k context length. The performance of ChunkKV is consistently better as the context length increases.
@@ B.4 Chunk Size
Table 15 shows the performance of ChunkKV with different chunk size on the LongBench benchmark.
Table 15: LongBench Performance Comparison with different chunk sizes
Model Chunk Size Full KV
3 5 10 20 30
LLaMA-3-8B-Instruct 40.49 40.47 40.51 40.05 39.57 41.46
Mistral-7B-Instruct 46.45 46.51 46.71 46.42 45.98 48.08
Qwen2-7B-Instruct 40.38 40.33 40.66 40.88 40.73 40.71
Table 16 shows the performance of ChunkKV with different chunk size on the GSM8K benchmark. Figure 19 shows that the ChunkKV with different chunk sizes on GSM8K displays the same curve pattern as LongBench. The CoT prompt length for GSM8K is only 1K tokens, so the optimal chunk size range is smaller.
@@ Appendix C Theoretical Understanding
Language Modeling. Current LLMs (Brown et al., 2020; Touvron et al., 2023; Fang & Xie, 2022) usually utilize the next word prediction as the language modelling, which predicts the next token $o_{t}$ given the previous tokens $o_{1}\dots o_{t-1}$ for all $t=1,\dots,T$. Formally, a language modelling can be writen as the distribution $f(o_{t}|o_{1:t-1})$. And it is pretrained on a huge corpus sampled from the pretraining distribution $p(o_{1},\dots,o_{t+1})$ (Fang & Xie, 2022). Considering the large scale of the model size and dataset size, it can be assumed that the $f(o_{1}\dots o_{t+1})$ has been aligned with the $p(o_{1}\dots o_{t+1})$ (Fang & Xie, 2022).
Prompt Distribution. Following (Fang & Xie, 2022), a prompt is composed of an input token sequence $x$ followed by an output token $y$. Then, the $i$-th training example  [^1] Here, training example in prompts means happens during the prompt learning. that can appear in any place in the whole prompt $o_{1:T}$ is defined as $O_{i}$ consisting of an input $x_{i}=O_{i}\left[1:k-1\right]$ (the first $k-1$ tokens) followed by the output $y_{i}=O_{i}\left[k\right]$ at the end, where the length $k$ is fixed for simplicity.
@@
In-context learning setups and Assumptions. We follow other settings and assumptions in  (Fang & Xie, 2022). With the greedy decoding (Fu et al., 2024b), sampling the next token from the language modeling $f(o_{t}|o_{1:t-1})$ becomes the predictor as $y=\operatorname*{arg\,max}_{o_{t}}f(o_{t}|o_{1:t-1})$.
@@
We define $p_{\theta}^{i}(o):=p(O[i]=o|O[1:i-1],\theta)$ of the $i$-th token with previous tokens and the analogous distribution $p^{i}_{prompt}:=p_{prompt}(O[i]=o|O[1:i-1])$ under the prompt distribution. Following (Fang & Xie, 2022), there is a distinguishability condition formalizes when in-context learning occurs giving the concept $\theta$.
@@
Noises from KV Cache Compression. Naturally, because of the sparsified KV cache, some history tokens in $o_{1:t-1}$ at different layers lost its attention score calculation with respect to the next word prediction $o_{t}$. We can regard this as the noise added onto the $o_{1:t-1}$. Thus, distincting $\theta^{\star}$ from $\theta$ requires larger KL divergence. Following (Zhou et al., 2024), we provide the following second condition about the distinguishability with the KV cache sparsity.
@@
Lemma 1 (noisy-relaxed bound in  (Fang & Xie, 2022; Zhou et al., 2024)).
@@
where $g(\nu)=\frac{1}{2}\big{(}(1-\nu)\log(1-\nu)+(1+\nu)\log(1+\nu)\big{)}$ is the calibration function (Steinwart, 2007; Pires & Szepesvári, 2016) for the multiclass logistic loss for $\nu\in[0,1]$.
Following (Kleijn & der Vaart, 2012; Fang & Xie, 2022), KL divergence is assumed to haver the 2nd-order Taylor expansion with the concept $\theta$. Then, we have the following theorem and proof.
@@
(Fang & Xie, 2022; Zhou et al., 2024) Let the set of $\theta$ which does not satisfy Equation 3 in Condition 1 to be $\mathcal{B}$. Assume that KL divergences have a 2nd-order Taylor expansion around $\theta^{\star}$:
@@
(Fang & Xie, 2022) By the Taylor expansion on $\theta$, we have for any $\theta$ in $\mathcal{B}$ that
@@
Token-level Importance Measure. The token-level KV cache methods usually calculate the importance of different tokens. Then, the KV cache with indexes that have higher importance will be preserved. Such indexes are normaly choosed as the attention score. Considering that the case in Figure 1, where each token in the $i$-th training  [^2] Here, training means prompt learning (Fang & Xie, 2022). example sequence ($O_{i}=[x_{i},y_{i}]$) might be compressed, and tokens are sparsified concretely without out dependency to other tokens. However, in the generation process of the $i$-th training example, $O_{i}=[x_{i},y_{i}]$ is sampled from $p(O_{i}|h^{\text{start}}_{i},\theta^{\star})$ and $p_{\theta}^{j}(o):=p(O[j]=o|O[1:j-1],\theta)$ of the $j$-th token with previous tokens and the analogous distribution $p^{j}_{prompt}:=p_{prompt}(O[j]=o|O[1:j-1])$. And the KL divergence is defined as $\text{KL}_{j}(\theta^{\star}||\theta):=\mathbb{E}_{O[1:j-1]\sim p_{prompt}}[\text{KL}(p^{j}_{prompt}||p_{\theta}^{j})]$, which means that in a training example $O_{i}=[x_{i},y_{i}]=O_{i}[1:k]$, each token $O_{i}[j]$ has strong dependency with $O_{i}[1:j-1]$, noises on previous any $j$-th token will influence the distinguishability on the following tokens, i.e. requiring larger $\left\{\text{KL}_{u}(\theta^{\star}||\theta)\right\}_{u>j}$.
@@ Appendix D Additional Related Work
KV cache sharing Recent work has explored various strategies for sharing KV caches across transformer layers. Layer-Condensed KV Cache (LCKV) (Wu & Tu, 2024) computes KVs only for the top layer and pairs them with queries from all layers, while optionally retaining standard attention for a few top and bottom layers to mitigate performance degradation. Similarly, You Only Cache Once (YOCO) (Sun et al., 2024) computes KVs exclusively for the top layer but pairs them with queries from only the top half of layers, employing efficient attention in the bottom layers to maintain a constant cache size. In contrast, Cross-Layer Attention (CLA) (Brandon et al., 2024) divides layers into groups, pairing queries from all layers in each group with KVs from that group’s bottom layer. MiniCache (Liu et al., 2024a) introduces a novel method that merges layer-wise KV caches while enabling recovery during compute-in-place operations, optimizing KV cache size. These methods illustrate various trade-offs between computation, memory usage, and model performance when sharing KV caches across transformer layers.
@@
Long-Context Benchmarks The landscape of long-context model benchmarks has evolved to encompass a wide range of tasks, with particular emphasis on retrieval and comprehension capabilities. Benchmarks for understanding have made significant strides, with $\infty$-Bench (Zhang et al., 2024) pushing the boundaries by presenting challenges that involve more than 100,000 tokens. LongBench (Bai et al., 2024) has introduced bilingual evaluations, addressing tasks such as long-document question answering, summarization, and code completion. Complementing these efforts, ZeroSCROLLS (Shaham et al., 2023) and L-Eval (An et al., 2023) have broadened the scope to include a diverse array of practical natural language tasks, including query-driven summarization.
@@
In parallel, retrieval benchmarks have largely relied on synthetic datasets, offering researchers precise control over variables such as the length of input tokens. This approach minimizes the impact of disparate parametric knowledge resulting from varied training methodologies. A significant body of recent work has concentrated on the development of synthetic tasks specifically for retrieval evaluation (Kamradt, 2023; Mohtashami & Jaggi, 2023; Li et al., 2023; Liu et al., 2024c; Hsieh et al., 2024). In addition, researchers have explored the potential of extended contexts in facilitating various forms of reasoning (Tay et al., 2021).
@@
This dual focus on synthetic retrieval tasks and comprehensive understanding benchmarks reflects the field’s commitment to rigorously assessing the capabilities of long-context models across diverse linguistic challenges. Prompting Compression In the field of prompt compression, various designs effectively combine semantic information to compress natural language. Wingate et al. (2022) utilize soft prompts to encode more information with fewer tokens. Chevalier et al. (2023) present AutoCompressor, which uses soft prompts to compress the input sequence and extend the original length of the base model. Both Zhou et al. (2023) and Wang et al. (2023) recurrently apply LLMs to summarize input texts, maintaining long short-term memory for specific purposes such as story writing and dialogue generation. The LLMLingua series (Jiang et al., 2023b, 2024; Fei et al., 2024) explores the potential of compressing LLM prompts in long-context, reasoning, and RAG scenarios. Fei et al. (2024) use pre-trained language models to chunk the long context and summarize semantic information, compressing the original context.
@@ Appendix F Statistics of Datasets
LongBench is a meticulously designed benchmark suite that evaluates the capabilities of language models in handling extended documents and complex information sequences. This benchmark was created for multi-task evaluation of long-context inputs and includes 17 datasets covering tasks such as single-document QA (Kočiský et al., 2018; Dasigi et al., 2021), multi-document QA (Yang et al., 2018; Ho et al., 2020; Trivedi et al., 2022; He et al., 2018), summarization (Huang et al., 2021; Zhong et al., 2021; Fabbri et al., 2019; Wu et al., 2023), few-shot learning (Li & Roth, 2002; Gliwa et al., 2019; Joshi et al., 2017), synthetic tasks and code generation (Guo et al., 2023; Liu et al., 2024d). The datasets feature an average input length ranging from $1K$ to $18K$ tokens, requiring substantial memory for KV cache management.
@@
GSM8K (Cobbe et al., 2021) 7,473 1,319
LongBench (Bai et al., 2024) - 4,750
NIAH* (Kamradt, 2023) - 800

### v3

@@ B.2 LongBench
Table 17 shows additional experiments with LLaMA-3-70B-Instruct, comparing our ChunkKV method with baselines with KV cache size = 256 on the LongBench dataset:
Table 17: Comprehensive performance comparison of ChunkKV across LongBench English subtasks with 70B model.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
LlaMa-3-70B-Instruct, KV Size = Full
FullKV 27.56 46.00 49.75 50.86 56.16 29.77 31.66 21.80 27.42 75.50 92.58 46.08 12.00 73.50 45.02 68.74 47.15
LlaMa-3-70B-Instruct, KV Cache Size = 256
H2O 26.42 37.52 36.23 47.82 49.34 24.30 26.96 20.67 26.50 72.00 92.16 43.56 12.00 73.00 47.21 63.13 43.67
SnapKV 26.79 42.23 48.78 50.89 55.46 28.14 23.85 22.20 24.62 73.00 93.05 44.53 12.00 72.50 54.24 65.60 46.11
ChunkKV 27.37 45.95 49.23 50.51 55.49 30.28 30.47 22.63 26.30 75.50 92.58 46.51 12.00 73.50 49.51 68.05 47.24
@@ B.3 Needle-In-A-Haystack
Figure 16 and 17 visualizes the performance of ChunkKV on the NIAH benchmark for LLaMA-3-8B-Instruct and Mistral-7B-Instruct with a KV cache size of 128 under 8k and 32k context length. The performance of ChunkKV is consistently better as the context length increases.
@@ B.4 Chunk Size
Table 18 and 19 show the performance of ChunkKV with different comperession ratios and different chunk sizes on the LongBench and NIAH. We conducted extensive experiments across different compression ratios and KV cache sizes to shows the effectiveness of ChunkKV and the chunk size is robust.
Compression Chunk Size
Rate 1 3 5 10 15 20 30
10% 37.32 40.49 40.47 40.51 40.21 40.05 39.57
20% 38.80 40.66 40.57 40.74 40.53 40.46 40.04
30% 39.23 41.02 41.29 41.59 41.38 41.33 41.02
KV Cache Chunk Size
Size 1 3 5 10 15 20 30
96 41.0 63.2 65.2 70.3 67.2 65.3 53.1
128 47.9 65.6 69.1 73.8 72.3 72.0 71.2
256 61.7 70.3 71.2 74.1 73.2 72.3 71.1
512 68.6 72.6 72.5 74.5 74.3 74.0 72.6
Figure 18 shows the performance of the ChunkKV with different chunk size on the LongBench and NIAH benchmarks. The three colorful curves represent three LLMs with different chunk sizes, and the colorful dashed line is the corresponding FullKV performance. From Figure 18, we can observe that the LongBench performance of ChunkKV is not significantly affected by the chunk size, with performance variations less than 1%. The three curves are closely aligned, indicating that chunk sizes in the range of $\{10,20\}$ exhibit better performance.
Table 20 and Figure 19 shows that the ChunkKV with different chunk sizes on GSM8K displays the same curve pattern as LongBench. The CoT prompt length for GSM8K is only 1K tokens, so the optimal chunk size range is smaller.
@@ B.5 Multi-Lingual
B.6 KV Cache Quantization
For comprehensively evaluate the effectiveness of ChunkKV, we conducted experiments comparing ChunkKV with quantization methods KIVI [37]. While both approaches aim to optimize LLM inference, they operate on fundamentally different principles: quantization reduces KV matrix precision, whereas our eviction method reduces KV matrix size.
From an implementation perspective, quantization methods require the full KV cache during prefilling to produce quantized representations, which are then used during decoding. In contrast, ChunkKV employs token removal prior to prefilling, enabling operation with a compressed cache throughout the entire inference process. This distinction creates different efficiency profiles, and each method offers unique advantages.
The following tables present our detailed analysis comparing ChunkKV with KIVI, across both performance and efficiency metrics on the LLaMA-3-8B-Instruct model. It should be noted that due to KIVI’s dependency on older Python versions, there may be some discrepancies between the efficiency results reported here and those in Table 8 of our paper.
Table 22: Comprehensive performance comparison of ChunkKV and quantization methods across LongBench English subtasks.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
LlaMa-3-8B-Instruct, KV Size = Full
ChunkKV with different compression ratios
ChunkKV-10% 24.89 22.96 37.64 43.27 36.45 20.65 22.80 22.97 20.82 71.50 90.52 40.83 5.93 69.00 60.49 57.48 40.51
ChunkKV-20% 26.13 28.43 38.59 44.46 34.13 21.06 24.72 23.11 22.91 71.50 90.56 41.51 5.09 69.00 58.17 52.51 40.74
ChunkKV-30% 25.88 29.58 38.99 43.94 34.16 21.70 26.50 23.15 23.95 72.00 90.56 42.47 5.34 69.25 61.68 56.35 41.59
KIVI with different quantization nbits
KIVI-2bits 25.45 29.45 40.85 45.25 35.55 22.05 25.48 22.95 26.06 70.25 89.00 41.73 4.25 64.00 59.72 46.70 40.54
KIVI-4bits 25.57 29.60 41.00 45.40 35.70 22.20 28.78 23.00 26.49 72.50 89.84 42.16 4.35 64.52 59.91 46.84 41.11
KIVI-8bits 25.70 29.73 41.11 45.52 35.81 22.34 25.62 23.01 26.19 72.97 90.54 41.83 4.62 69.22 58.02 50.74 41.43
Table 22 demonstrates that ChunkKV with 30% compression ratio achieves comparable performance (41.59) to KIVI with 8-bit quantization (41.43) across LongBench subtasks. Notably, ChunkKV shows particularly strong results in code-related tasks, while KIVI maintains better performance in single-document QA scenarios.
These results highlight the complementary strengths of eviction and quantization approaches. ChunkKV’s semantic chunk-based method delivers particularly strong efficiency benefits while maintaining competitive performance with state-of-the-art quantization techniques.
B.7 Efficiency results
Table 23 shows the efficiency results for different KV cache strategies with varying output lengths and the metrics are Time to First Token (TTFT) and Token Processing Time (TPOT).
Table 23: Efficiency Results for Different KV Cache Strategies with Varying Output Lengths
Method Input Output TTFT(s) $\downarrow$ TPOT(ms/token) $\downarrow$
Output Length = 1024
FullKV 4096 1024 1.17 42.58
StreamingLLM 4096 1024 1.02 (12.8%) 35.94 (15.6%)
SnapKV 4096 1024 1.04 (11.1%) 36.24 (14.9%)
ChunkKV 4096 1024 1.03 (12.0%) 36.64 (13.9%)
ChunkKV_reuse 4096 1024 1.00 (14.5%) 36.47 (14.3%)
Output Length = 4096
FullKV 4096 4096 1.17 42.85
StreamingLLM 4096 4096 1.09 (6.8%) 40.01 (6.6%)
SnapKV 4096 4096 1.10 (6.0%) 40.24 (6.1%)
ChunkKV 4096 4096 1.10 (6.0%) 40.17 (6.3%)
ChunkKV_reuse 4096 4096 1.08 (7.7%) 40.16 (6.3%)
@@ Appendix C Theoretical Understanding
Language Modeling. Current LLMs [1, 3, 20] usually utilize the next word prediction as the language modelling, which predicts the next token $o_{t}$ given the previous tokens $o_{1}\dots o_{t-1}$ for all $t=1,\dots,T$. Formally, a language modelling can be writen as the distribution $f(o_{t}|o_{1:t-1})$. And it is pretrained on a huge corpus sampled from the pretraining distribution $p(o_{1},\dots,o_{t+1})$ [20]. Considering the large scale of the model size and dataset size, it can be assumed that the $f(o_{1}\dots o_{t+1})$ has been aligned with the $p(o_{1}\dots o_{t+1})$ [20].
[…]
@@
In-context learning setups and Assumptions. We follow other settings and assumptions in  [20]. With the greedy decoding [38], sampling the next token from the language modeling $f(o_{t}|o_{1:t-1})$ becomes the predictor as $y=\operatorname*{arg\,max}_{o_{t}}f(o_{t}|o_{1:t-1})$.
@@
We define $p_{\theta}^{i}(o):=p(O[i]=o|O[1:i-1],\theta)$ of the $i$-th token with previous tokens and the analogous distribution $p^{i}_{prompt}:=p_{prompt}(O[i]=o|O[1:i-1])$ under the prompt distribution. Following [20], there is a distinguishability condition formalizes when in-context learning occurs giving the concept $\theta$.
@@
Noises from KV Cache Compression. Naturally, because of the sparsified KV cache, some history tokens in $o_{1:t-1}$ at different layers lost its attention score calculation with respect to the next word prediction $o_{t}$. We can regard this as the noise added onto the $o_{1:t-1}$. Thus, distincting $\theta^{\star}$ from $\theta$ requires larger KL divergence. Following [39], we provide the following second condition about the distinguishability with the KV cache sparsity.
@@
Lemma 1 (noisy-relaxed bound in  [20, 39]).
@@
where $g(\nu)=\frac{1}{2}\big{(}(1-\nu)\log(1-\nu)+(1+\nu)\log(1+\nu)\big{)}$ is the calibration function [40, 41] for the multiclass logistic loss for $\nu\in[0,1]$.
Following [42, 20], KL divergence is assumed to haver the 2nd-order Taylor expansion with the concept $\theta$. Then, we have the following theorem and proof.
@@
[20, 39] Let the set of $\theta$ which does not satisfy Equation 3 in Condition 1 to be $\mathcal{B}$. Assume that KL divergences have a 2nd-order Taylor expansion around $\theta^{\star}$:
@@
[20] By the Taylor expansion on $\theta$, we have for any $\theta$ in $\mathcal{B}$ that
@@
[…]
@@ Appendix D Additional Related Work
Chunking Method. The chunking methodology is widely used in the field of NLP due to its simplicity and effectiveness [43]. In the era of LLMs, chunking is primarily applied in data pre-processing. For example, Shi et al. [44] suggest grouping related training data into chunks to achieve better convergence curves to pre-train LLMs. Fei et al. [45] apply a topic-based chunking method to improve the semantic compression of prompts. Furthermore, chunking plays an important role in the Retrieval-Augmented Generation (RAG) field [46, 47, 48]. It serves to divide documents into units of information with semantic content suitable for embedding-based retrieval and processing by LLMs.
Layer-Wise Technique. The layer-wise technique is widely used in the training and inference of large language models (LLMs). LISA [49] is a layer-wise sampling method based on observations of the training dynamics of Low-Rank Adaptation (LoRA)[50] across layers. LAMB[51] is a layer-wise adaptive learning rate method that speeds up LLM training by stabilizing training convergence with large batch sizes. DoLa [52] employs layer-wise contrasting to reduce hallucinations during LLM inference.
@@
KV Cache Sharing. Recent work has explored various strategies for sharing KV caches across transformer layers. Layer-Condensed KV Cache (LCKV) [53] computes KVs only for the top layer and pairs them with queries from all layers, while optionally retaining standard attention for a few top and bottom layers to mitigate performance degradation. Similarly, You Only Cache Once (YOCO) [54] computes KVs exclusively for the top layer but pairs them with queries from only the top half of layers, employing efficient attention in the bottom layers to maintain a constant cache size. In contrast, Cross-Layer Attention (CLA) [55] divides layers into groups, pairing queries from all layers in each group with KVs from that group’s bottom layer. MiniCache [56] introduces a novel method that merges layer-wise KV caches while enabling recovery during compute-in-place operations, optimizing KV cache size. These methods illustrate various trade-offs between computation, memory usage, and model performance when sharing KV caches across transformer layers.
@@
Long-Context Benchmarks. The landscape of long-context model benchmarks has evolved to encompass a wide range of tasks, with particular emphasis on retrieval and comprehension capabilities. Benchmarks for understanding have made significant strides, with $\infty$-Bench [57] pushing the boundaries by presenting challenges that involve more than 100,000 tokens. LongBench [21] has introduced bilingual evaluations, addressing tasks such as long-document question answering, summarization, and code completion. Complementing these efforts, ZeroSCROLLS [58] and L-Eval [59] have broadened the scope to include a diverse array of practical natural language tasks, including query-driven summarization.
@@
In parallel, retrieval benchmarks have largely relied on synthetic datasets, offering researchers precise control over variables such as the length of input tokens. This approach minimizes the impact of disparate parametric knowledge resulting from varied training methodologies. A significant body of recent work has concentrated on the development of synthetic tasks specifically for retrieval evaluation [22, 60, 61, 62, 63]. In addition, researchers have explored the potential of extended contexts in facilitating various forms of reasoning [64].
This dual focus on synthetic retrieval tasks and comprehensive understanding benchmarks reflects the field’s commitment to rigorously assessing the capabilities of long-context models across diverse linguistic challenges.
@@
Prompting Compression. In the field of prompt compression, various designs effectively combine semantic information to compress natural language. [65] utilize soft prompts to encode more information with fewer tokens. [66] present AutoCompressor, which uses soft prompts to compress the input sequence and extend the original length of the base model. Both [67] and [68] recurrently apply LLMs to summarize input texts, maintaining long short-term memory for specific purposes such as story writing and dialogue generation. The LLMLingua series [69, 70, 45] explores the potential of compressing LLM prompts in long-context, reasoning, and RAG scenarios. [45] use pre-trained language models to chunk the long context and summarize semantic information, compressing the original context.
KV Cache Quantization. Among quantization approaches for KV cache optimization, SmoothQuant [71] stands out as a notable post-training technique. By implementing balanced transformations between activation and weight quantization difficulties, this method enables 8-bit KV cache compression with minimal performance degradation. However, research by  [72] has demonstrated that SmoothQuant encounters substantial accuracy deterioration when compressed beyond 8 bits. In parallel developments, the FlexGen system [73] implements group-wise quantization at 4 bits for both key and value caches, while KIVI [37] introduces an innovative approach with asymmetric 2-bit per-channel quantization that requires no additional tuning phase.
@@ Appendix F Statistics of Datasets
LongBench is a meticulously designed benchmark suite that evaluates the capabilities of language models in handling extended documents and complex information sequences. This benchmark was created for multi-task evaluation of long-context inputs and includes 17 datasets covering tasks such as single-document QA [74, 75], multi-document QA [76, 77, 78, 79], summarization [80, 81, 82, 83], few-shot learning [84, 85, 86], synthetic tasks and code generation [87, 88]. The datasets feature an average input length ranging from $1K$ to $18K$ tokens, requiring substantial memory for KV cache management.
@@
GSM8K [23] 7,473 1,319
LongBench [21] - 4,750
NIAH* [22] - 800
JailBreakV [24] - 28,000